SDAS Users Guide

You will receive the SDAS package in compressed tar format (sdas.tar.Z). To extract the SDAS package from sdas.tar.Z simply cd to a convienient directory and type

  uncompress < sdas.tar.Z | tar -xvf -

This will create the necessary subdirectories and extract the appropriate files. The next step is to build the SDAS libraries and install the software. Note: If you want to use the default configuration then simply type make install at this point (you might need to be superuser to do this)

1. Edit Makefile to reflect your particular configuration:

(a) Change BINDIR = /usr/local/bin to reflect the subdirectory where the executable utilities for SDAS are to be located.

(b) Change LIBDIR = /usr/local/lib to reflect the subdirectory where the SDAS libraries will be located.

(c) Change SRCDIR = /usr/local/src to reflect the subdirectory where the files sdas.c and sdas.h will be located.

(d) Change NAME = sdas to reflect the name you want the executable SDAS program to be (also, this will be the basename of the PDT file and the primary PCT file).

(e) Change SBINDIR = . to reflect the directory where the executable SDAS program (sdas) will be located.

2. Type make install

The procedure outlined above will install SDAS on your system it will also produce a shell script that will be useful in the next step. The next step is to configure SDAS for a particular application.

1. Decide where the configured SDAS executable program will be located. The default is the current directory.

2. Create the necessary PDT and PCT files (NAME.pdt and NAME.pct).

3. Edit the configure file (shell script):

(a) Set ULIB = to reflect the path and name of the library of routines to be linked to the SDAS program.

(b) Set UOBJ = to reflect the names of the object files to be linked to the SDAS program.

(c) Set CDIR = . to reflect the directory where the configuration files (NAME.pct and NAME.pdt) are located.

(d) Change SBINDIR = . to reflect the directory where the executable SDAS program (sdas) will be located.

4. Type configure

This procedure will produce a SDAS program NAME in the current directory (or the directory specified by SBINDIR. To add additional routines the NAME simply include the routine’s name in the appropriate PCT file, add the name of the object file containing the routine to UOBJ, then type configure.